OpenStudy (anonymous):

please solve!! \[\left| 3x-9 \right|=6\]

OpenStudy (anonymous):

\[\pm (3x - 9) = 6\]

OpenStudy (anonymous):

Solve it just like the last one we did!

OpenStudy (anonymous):

3x - 9 = 6 3x = 15 x = 5 -3x + 9 = 6 -3x = -3 x = 1

OpenStudy (anonymous):

Thank you both :D